How many ounces of pure bran flakes must be added to 65 oz of cereal that is 21% bran flakes to produce a mixture that is 35% bran flakes?

Answer by richwmiller(17219)

We have 65 oz of 21% bran
We want to add some 100% bran to make 35% bran
1*x+0.21*65=0.35(65+x)
1*x+13.65=22.75+0.35x
1x-0.35x=22.75-13.65
0.65x=9.1
x=14 oz at 100% (14) is added to 65 oz at 21%(13.65) to make 35% of 79 oz(27.65)
check
1*14+0.21*65=0.35(65+14)
14+13.65=0.35(79)
27.65=27.65
ok
codemixadd